# A Survey on Transformer Compression

### 05 Feb 2024 v1

@@ 6 Conclusions and Future directions
In this survey, we systematically investigate the compression methods for transformer models. Compared with compression method for conventional models, there are unique considerations when compressing transformer models. Unlike other architectures such as CNN or RNN, the transformer a unique architecture design with alternative attention and FFN modules, necessitating specifically tailored compression methods for optimal compression rates. Moreover, the efficiency of compression methods becomes particularly crucial for these large models. Certain model compression techniques demand substantial computational resources, which may prove prohibitive for such extensive models.This survey aims to encompass the majority of recent works pertaining to transformers and articulate a comprehensive roadmap for their compression. Subsequently, we delve into the interconnections among various methods, address post-challenges, and outline directions for future research.
Relationship between different compression methods. Different compression methods can be used together to obtain an extremely efficient architecture. A conventional sequence is to firstly define a new architecture with efficient operations. The redundant components (e.g., attention head, layers) are then removed to get a smaller model. For practical hardware implementation, quantizing weights or activations to lower bits is imperative. The choice of required bits depend on not only the tolerance of error, but also hardware design. As an illustration, Int8 computation is efficiently optimized on Nvidia A00, but it lacks support on the older Tesla P100. Distillation commonly serves as a training strategy, applicable during the fine-tuning phases of both pruning and quantization. For pursing extremely high compression rate, it is promising to explore how combining different compression strategies. Though has been well explored in conventional models like CNN [9, 284], the transformer models have more complicated architectures and higher computational cost. It is challenging to find a proper combination strategy via joint search.
@@
Different from compression for conventional models, the computational cost for compression methods has been put on higher priority. Large transformers are currently trained on vast datasets using significant computational resources. For instance, Llama2 is trained on 2 trillion tokens with thousands of GPUs over several months [1]. It is impractical to fine-tune a model with comparable computational resources during pre-training, especially when the original data is often inaccessible. Therefore, the feasibility of efficient compression methods after training becomes more viable.
@@
Initially developed for conventional small models, a series of works have extensively researched post-training quantization [21, 41, 22, 42], and these methods have seamlessly transitioned to transformers. With only several GPU hours, some recent works GPTQ [73] ,SmoothQuant [18] quantizes a FP16 model to Int8 without significant performance loss. However, for lower bits (e.g., 4 bit), the quantized model still surfers significant performance degradation [73, 18]. It is worth noting that extremely-low-bit models, such as binary transformers, have been extensively explored in conventional small models but remain relatively unexplored in the context of large models.
@@
For pruning, the challenge of post-training is intricately linked to the pruning granularity. Though the nonstructural sparsity can achieve high compression rate with minimal fine-tuning requirements [143], the similar strategy is hard to transfer to the structural pruning. Directly removing the entire attention heads or layers will result in a substantial alteration of the model’s architecture and a consequent significant reduction in accuracy. How to recognize effective weights and how to effectively recover the performance are both insight directions. The efficient strategy to identify effective weights and recovery representation ability are key research directions in addressing these challenges.
@@
Efficient architectures beyond Transformer. In real-world applications, the input context for a transformer architecture can extend to an extremely long length, encompassing sequence texts (e.g., a book with hundreds of thousands of words) in NLP or high-resolution images in CV. The vanilla attention mechanism exhibits quadratic complexity w.r.t. the length of the input sequence, posing a significant computational challenge for long-sequence inputs. Numerous studies have addressed this issue by mitigating the computational cost of attention, employing techniques such as sparse attention, local attention, etc(See Section 4.2). However, these attention compression strategies often compromise the representation ability, leading to diminished performance.
@@
Emerging architectures such as RWKV [216] and RetNet [36] adopt a recursive output generation akin to RNNs, effectively reducing computational complexity to $\mathcal{O}(N)$. This development holds promise for further exploration in the quest for more efficient models. For computer vision tasks, even a pure MLP architecture without attention module can achieve SOTA performance [40, 241, 257, 255, 256]. Beyond the widely used Transformer architecture, it is promising to explore new efficient architecture by carefully investigating their efficiency, generalization and scaling ability.

### 07 Apr 2024 v2

@@ 6 Conclusions and Future Directions
In this survey, we systematically investigate the compression methods for Transformer models. Compared with compression method for conventional models, there are unique considerations when compressing Transformer models. Unlike other architectures such as CNN or RNN, the Transformer adopts a unique architecture design with alternative attention and FFN modules. This necessitates specifically tailored compression methods for optimal compression rates. Moreover, the efficiency of compression methods becomes particularly crucial for large models. Certain model compression methods require substantial amounts of computational resources, making them potentially prohibitive for such models. This survey aims to encompass the majority of recent works pertaining to Transformers and articulate a comprehensive roadmap for their compression. Subsequently, we delve into the interconnections among various methods, address post-challenges, and outline directions for future research.
Relationship between different compression methods. Different compression methods can be used together to obtain an extremely efficient architecture. A conventional sequence is to firstly define a new architecture with efficient operations. Any redundant components (e.g., attention head, layers) are then removed to reduce the model size. For practical hardware implementation, quantizing weights or activations to lower bits is imperative. The choice of required bits depend on not only the tolerance of error, but also hardware design. As an illustration, Int8 computation is efficiently optimized on Nvidia A100 GPU, but it lacks support on the older Tesla P100 GPU. Distillation commonly serves as a training strategy that applies during the fine-tuning phases of both pruning and quantization. Combining different compression strategies is a promising direction to explore for achieving extremely high compression rates. Although detailed exploration has been conducted into conventional models like CNN [9, 289], Transformer models have more complicated architectures and higher computational costs. Finding a suitable combination strategy via joint search is a challenging task.
@@
Different from compression for conventional models, the cost of compressing large Transformers has been put on higher priority. Large Transformers are currently trained on vast datasets using significant computational resources. For instance, Llama2 is trained on 2 trillion tokens with thousands of GPUs over several months [1]. It is impractical to fine-tune a model with comparable computational resources during pre-training, especially when the original data is often inaccessible. Therefore, the feasibility of efficient compression methods after training becomes more viable.
@@
Initially developed for conventional small models, a series of works have extensively researched post-training quantization [21, 41, 22, 42], and these methods have seamlessly transitioned to Transformers. With only several GPU hours, some recent works GPTQ [72],and SmoothQuant [18] quantize an FP16 model to Int8 without significant performance loss. However, for lower bits (e.g., 4 bit), the quantized model still surfers significant performance degradation [72, 18]. It is worth noting that extremely-low-bit models, such as binary Transformers, have been extensively explored in conventional small models, yet they remain relatively unexplored in the context of large models.
@@
For pruning, the challenge of post-training is intricately linked to the pruning granularity. Although nonstructural sparsity can achieve a high compression rate with minimal fine-tuning requirements [143], a similar strategy is hard to transfer to the structural pruning. Directly removing the entire attention heads or layers will result in a substantial alteration of the model’s architecture and a significant reduction in accuracy. How to recognize effective weights and how to effectively recover the performance are both important directions for exploration. The efficient strategy to identify effective weights and recovery representation ability are key research directions in addressing these challenges.
@@
Efficient architectures beyond Transformer. In real-world applications, the input context for a Transformer architecture can be extremely long, encompassing sequence texts (e.g., a book with hundreds of thousands of words) in NLP or high-resolution images in CV. The vanilla attention mechanism exhibits quadratic complexity regarding the length of the input sequence, posing a significant computational challenge for long-sequence inputs. Numerous studies have addressed this issue by mitigating the computational cost of attention, employing techniques such as sparse attention, local attention, etc(See Section 4.2). However, these attention compression strategies often compromise the representation ability, leading to diminished performance.
@@
Emerging architectures such as RWKV [217] and RetNet [36] adopt a recursive output generation akin to RNNs, effectively reducing computational complexity to $\mathcal{O}(N)$. This development holds promise for further exploration in the quest for more efficient models. For computer vision tasks, even a pure MLP architecture without an attention module can achieve SOTA performance [40, 244, 260, 258, 259]. Beyond the widely used Transformer architecture, it is promising to explore new efficient architectures by carefully investigating their efficiency, generalization and scalability.